OpenStudy (anonymous):

Why is the series sigma n=1 to infinity of cos(n) divergent

OpenStudy (anonymous):

Because the terms do not approach zero.

OpenStudy (turingtest):

what happens as n increases? do the terms of the series\[\{\cos(n)\}_{n=1 }^\infty\]approach zero? no, they oscillate forever between 1 and -1, which means we are always adding more and more non-zero terms that do not vanish as \(n\to\infty\), hence the series diverges

OpenStudy (anonymous):

\(\cos \theta\) with n being integers: 

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \cos \theta\]